Periodic inventory by three methods; cost of goods sold

The units of an item available for sale during the year were as follows:

Jan. 1 Inventory 40 units at $126

Mar. 10 Purchase 40 units at $138

Aug. 30 Purchase 20 units at $142

Dec. 12 Purchase 100 units at $146

There are 80 units of the item in the physical inventory at December 31. The periodic inventory system is used.

Determine the ending inventory cost and the cost of goods sold by three methods. Round interim calculations to one decimal and final answers to the nearest whole dollar.

Cost of Ending Inventory and Cost of Goods Sold

  1. Inventory Method Ending Inventory Cost of Goods Sold
    The method of inventory costing based on the assumption that the cost of goods sold should be charged against revenue in the order in which the costs were incurred.First-in, first-out (FIFO) $ $
    A method of inventory costing based on the assumption that the most recent inventory costs should be charged against revenue.Last-in, first-out (LIFO)
    A method of inventory costing in which the cost of the units sold and in ending inventory is a weighted average of the purchase costs.Weighted average cost
Answer #1

* units of closing inventory = 80 units

* units of inventory sold = opening units + purchased units - closing units

= 40 + (40+20+100) -80

= 120 units

* CALCULATION OF COST OF GOODS SOLD AND COST OF CLOSING INVENTORY

1. By FIFO method:

cost of goods sold = (40 x 126) + (40 x 138) + (20 x 142) + (20 x 146) = 16320

Cost of closing inventory = 80 x 146 = 11680

2. By LIFO method:

Cost of goods sold = (100 x 146) + (20 X 142) = 17440

Cost of closing inventory = (40 X 138) + (40 X 126) = 10560

3. By Weighted Average method:

= 28000/200

= 140 per unit

Cost of goods sold = 120 x 140 = 16800

Cost of closing inventory = 80 x 140 = 11200

Inventory method ending inventory cost of goods sold
FIFO $11680 $16320
LIFO $10560 $17440
WEIGHTED AVERAGE METHOD $11200 $16800